# Which Augmentation Should I Use? An Empirical Investigation of Augmentations for Self-Supervised Phonocardiogram Representation Learning

### v3

@@ 2.3 Model Robustness in 1D biosignal classification
Extracting generalized and robust representations has been one of the most important long-standing goals of machine learning 6472238 . To test and improve the generalizabilty prowess and robustness of deep learning models, several fields have emerged, such as Domain Adaptation wang2018deep , Transfer Learning zhuang2020comprehensive , and Domain Generalization zhou2022domain ; 10233054 ; ballas_multi , which use data from non-i.i.d. distributions or out-of-distribution (OOD) data for model evaluation. These distinct data distributions are referred to as data Domains, as they are most often the outcome of different data generating processes. Here, we reference prior works which fall under the above research fields.
@@
Domain adaptation (DA) methods are the most popular choice when aiming to produce models with improved generalization capabilities in 1D biosignal classification. Similar to Transfer Learning (TL), DA algorithms leverage the representation learning capabilities of models pretrained on diverse source data distributions and fine-tune them on previously unseen but similar target distributions. The end-result is a model which remains robust under the distribution shift between source and target data. In healthcare, DA methods are proposed for a plethora of biosignal classification tasks. In weimann_transfer_2021 , the authors propose pre-training vanilla CNN networks on a large corpus of raw ECG signals and demonstrate their effectiveness when fine-tuned on a smaller dataset for arrhythmia detection. In another more recent work, He et al. he_novel_2023 propose tackling the distribution shift between different arrhythmia detection ECG datasets, by employing three distribution alignment mechanisms. To be more specific, they first extract spatio-temporal features from preprocessed signals, pass them through a graph convolutional neural network and finally align the resulting representations between source and target domains. Many studies have researched the advantages of DA and TL in EEG classification as well wan_review_2021 . To name a few, the authors of 10035017 ; 9154600 and peng_domain_2022 propose aligning source and target distributions via adversarial learning for motor image EEG classification and epilepsy detection respectively, while lan_domain_2018 investigates several state-of-the-art DA algorithms for EEG-based emotion recognition.
@@
Finally, Domain Generalization algorithms aim to tackle the model generalization at its core, as the main goal is to provide generalized models that perform well across unseen datasets. Although limited, there has been a recent interest in biosignal DG research. In the field of PCG classification, the authors of 9298838 propose a ensemble classifier fusion method yields improved results in a DG setup for abnormal vs normal signal classification. To tackle the domain shift and variability present in ECG signals collected from different patients, 9344210 and ma_reducing_2019 propose domain-adversarial model training for ECG and EEG classification, respectively. Most recently, the authors of 10233054 introduce a DG benchmark and a model which leverages representations from multiple layers of the network, inspired from 9898255 , for 12-lead ECG and 64-channel EEG classification.
@@ 3 Materials and Methods
We introduce an evaluation framework for assessing both the combination of various transformations and the robustness of a network trained via contrastive SSL, against a fully supervised baseline model. This allows us to evaluate and report on the most promising signal transformations or augmentations, for the PCG classification domain, while also highlight additional benefits of training an end-to-end classification model via Self-Supervised Contrastive Learning, such as improved model robustness and generalization.
@@
For the self-supervised training of all models, we adapt the SimCLR chen_simple_2020 methodology for 1D PCG representation learning and then evaluate numerous augmentation combinations designed specifically for 1D signal processing tasks.
@@
In the following sections, we describe the datasets and preprocessing steps followed for all signals included in the study, and describe in detail the protocol followed for evaluating the significance of each combination of augmentations, as well as the generalization ability of a model pretrained with self-supervised contrastive learning. Finally, we present the architecture and relevant implementation details of the 1D CNN network implemented for PCG classification.
3.1 Datasets & Downstream tasks
@@
The PCG signals used in this study originate from a total of 5 distinct public datasets and are available as “wav” files. These include the Ephnogram kazemnejad2021ephnogram , FPCGDB cesarelli_simulation_2012 , Pascal pascal-chsc-2011 , PhysioNet2016 liu_open_2016 , and PhysioNet2022 reyna_heart_2023 databases. Here we describe each dataset separately and provide all necesary details regarding the preprocessing steps followed for the homogenization of all signals into a common format.
@@
The Ephnogram kazemnejad2021ephnogram database contains a total of 69 multi-channel, simultaneous ECG and PCG recordings from 24 healthy adults aged between 23 and 29, acquired during physical activity and rest conditions. Most recordings (62 during exercise) are 30min long, while the duration of the remaining 7 (rest) is about 30 seconds. As specified by the researchers, the dataset includes poor sample signals for noise research purposes. As this dataset was proposed and gathered for multi-modal signal processing purposes and not for pathology detection or classification, it does not include labeled samples. For our study, we use the unlabeled samples of the PCG signals for the SSL pretraining of our models. All PCG recordings were sampled at 8 kHz.
@@
The FPCGDB cesarelli_simulation_2012 or Fetal PCG Database, as its name suggests, is a dataset containing 26 20-minute fetal phonocardiographic signals, gathered from healthy women during their the remaining months of their pregnancy and sampled at 333 Hz. Similarly to the Ephnogram dataset, the collected signals were obtained for the development of fetal PCG signal simulation and fetal heart rate extraction algorithms. Hence, this dataset also does not contain any labeled samples for classification purposes. Furthermore, most recordings are somewhat corrupted by noise, either due to the pregnant human body which contains various noise sources, such as muscular movements or placental blood turbulence, or even due to the surrounding environment.
@@
The Pascal pascal-chsc-2011 Heart Sound Classification challenge was introduced for developing algorithms for the automatic screening of cardiac pathologies. The challenges contains two tasks; one for heart sound segmentation and one for classification. In our work, we use the publicly available labeled data for model training and evaluation which was gathered in both clinical and in-the-wild settings. Notably, the 832 PCG signals are classified as either Normal, Murmur, Extra Heart Sound, Artifact or Extrasystole [^2] Each different category is described in the official challenge description (https://istethoscope.peterjbentley.com/heartchallenge/index.html).. The recording lengths vary from 1 to 30 seconds and are sampled at 195 Hz. As each sample is annotated with a single label, the above task is a multi-class classification problem.
@@
The remaining two datasets were also introduced in the context of public challenges for PCG classification. Similarly to the Pascal challenge, the objective of the PhysioNet2016 liu_open_2016 challenge aims at urging researchers to develop robust algorithms for differentiating normal and abnormal heart sound recordings but instead introduces a binary classification problem. The data included for model training in the dataset contains a total of 3,126 PCG recordings collected from a single precordial location from both healthy subjects and patients suffering from a variety of heart defections, lasting from 5 to around 120 seconds. Once again, the recordings were gathered in both clinical and non-clinical environments, with a sample rate of 2 kHz.
@@
Finally, the PhysioNet2022 challenge reyna_heart_2023 was designed for the development of algorithms which could detect heart murmur in PCG signals. Succedent to the 2016 challenge, the 2022 challenge contains multiple heart sound recordings, collected from multiple auscultation locations. The available data includes 3,163 recordings from 942 patients in total. Each patient has one or more PCG signals, recorded from one of the four heart valves or from a separate prominent auscultation location. Each of the recordings can be classified as ”present”, ”absent” or ”unknown”, meaning that heart murmur presence can be either detected, not detected or can not be determined (e.g due to noise or corrupted audio) respectively, rendering the challenge a multi-class classification problem. The sampling rate of the above signals is 4 kHz.
@@
Each dataset contains signals of varying length, sampling rate and label classes. To accommodate model training in our framework, we follow a similar process as in ballas_listen2yourheart_2022 for converting all signals into a similar format. As a first step, to avoid transient noise artifacts, we initially discard the first and final 2 seconds of each PCG recording. The resulting signal is split into 5 second overlapping windows with a 50% overlap. After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether it derives from a normal heart sound recording or not [^3] In this study, all recordings initially labeled as unknown are considered abnormal.. For the SSL pretraining stage, we assign pseudolabels to each window, as described in Section 3.2. The assignment of binary labels to each window enables the evaluation of trained models across datasets, which do not necessarily share similar sample classes. For example, a classifier trained to detect normal PCG signals in the Pascal dataset, can be then evaluated without further training on samples from the PhysioNet2022 dataset. The above evaluation process is described in detail in Section 3.4.
3.2 Self-Supervised Contrastive Learning for PCG Representation Learning
@@
The main goal of this study is to, first of all, verify the use of SSL contrastive learning and also evaluate the importance of augmentation selection, for robust PCG representation learning. The proposed framework consists of two separate tasks, a Pretraining pretext task and a Downstream task, as depicted in Figure 1.
@@
Assume a 1D PCG signal $x\in\mathbb{R}^{1\times L}$ sampled from $\mathcal{X}$, where $L$ is the length of the signal. The aim of the contrastive SSL pretext task is to learn robust representations of $x$ from unlabeled PCG signals, by maximizing the agreement between two different augmented versions of itself (positive pair) and simultaneously minimizing the agreement between entirely separate augmented windows. Essentially, a model is trained to predict whether two augmented signals are the result of different transformations applied on the original sample. Formally, let $t(\cdot)$ be a transformation or augmentation function drawn from a family of functions $\mathcal{T}$. During the pretraining phase, an audio window $x$ is augmented to two different versions $\tilde{x}$ and $\tilde{x}^{\prime}$, after the application of two transformation functions $t\sim T$ and $t^{\prime}\sim T^{\prime}$ respectively. Therefore, an initial batch of $N$ sample windows $(\tilde{x}_{1},\tilde{x}_{2},\dotsc,\tilde{x}_{n})$ results into an augmented batch of $2N$ signals $(\tilde{x}_{1},\tilde{x}_{2},\dotsc,\tilde{x}_{n},\tilde{x}_{1}^{\prime},\tilde{x}_{2}^{\prime},\dotsc,\tilde{x}_{n}^{\prime})$, where $\tilde{x}_{i}=t(\tilde{x})$ and $\tilde{x}_{i}^{\prime}=t^{\prime}(\tilde{x}_{i})$ for $i\in[1,N]$. As mentioned, each augmented pair of signals resulting from the same original window is considered a positive pair, e.g $(\tilde{x}_{1},\tilde{x}_{1}^{\prime})$, while pairs of different original windows comprise negative pairs, e.g $(\tilde{x}_{1},\tilde{x}_{2}^{\prime})$. Following its augmentation, the $2N$ signal batch is passed through a neural network encoder model $E(\cdot)$ (Figure 2), where each signal is encoded into a representation $\tilde{h}=E(\tilde{x})$. Subsequently, the resulting representations are flattened and passed through a fully connected linear layer $G(\cdot)$ and projected into 128-dimensional vectors $\tilde{z}=G(\tilde{h})$.
@@
For the contrastive prediction task, we employ the contrastive NT-Xent loss function introduced in oord_representation_2019 and implemented in the SimCLR framework chen2020simple . Specifically, given a set $\{\tilde{x}_{k}\}$ which includes an augmented positive pair $\tilde{x}_{i}$ and $\tilde{x}_{j}$, a model trained via contrastive SSL aims to identify $\tilde{x}_{j}$ in $\{\tilde{x}_{k}\}_{k}\neq i$ for a given $\tilde{x}_{i}$. The contrastive loss for a positive pair $(\tilde{x}_{i},\tilde{x}_{j})$ is formulated as follows:
@@
3.3 Augmentations for PCG Representation Learning
@@ 3.4 Evaluation Protocol
In the first step of the framework, we create a pool of 5 sec unlabeled audio windows from all available datasets, namely Pascal, PhysioNet2016, PhysioNet2022, Ephnogram and Fpcgdb. After the preprocessing step has been completed and all signals are homogenized in the common format described in Section 3.1, they are randomly shuffled and split into training and validation splits (80% and 20% respectively). In the second step, the split signals are cut into batches. Each sample in the batch is then duplicated and augmented by applying either one or a combination of the functions described in Section 3.3. The augmented samples are then passed through a Neural Network encoder, which is trained via the contrastive pretraining task (Step 3). After the SSL backbone model has been trained, its weights are frozen.
@@
In Step 4, a classification head with learnable weights is attached to the pretrained encoder which is then trained on data from one of the labeled datasets (Pascal, PhysioNet2016 or PhysioNet2022). Specifically, each of these datasets is split into training, validation and test splits with a standard 70%-20%-10% ratio and the trained model is evaluated on the test split only. Since each model is evaluated on test data originating from the same distribution as its training data, we call this the ”In-Distribution” evaluation. Inspired by recent work in biosignal domain generalization 10233054 , we also design an evaluation protocol for assessing the robustness of a trained classifier on out-of-distribution data, which are drawn from entirely different datasets. Since the data has been completely left-out during downstream training, the above process is followed for evaluating the generalization ability of a model. Specifically, we now also freeze the weights of the attached classification layers and test the model on data drawn from the remaining 2 labeled datasets. For example, if the network is trained for the Pascal downstream task, it is evaluated on both in-distribution data (Pascal) and on out-of-distribution data (PhysioNet2016 and PhysioNet2022). We call this step the out-of-distribution or OOD evaluation step. However, in order to evaluate the same model on each separate dataset, all data must share common labels. To that end, we implement two different downstream task types: ‘All’ and ‘Binary‘. In the ‘All‘ downstream task, the original labels of each dataset are assigned to each audio window, while in the ‘Binary’ task a signal is labeled as either ‘normal’ or ‘abnormal’, as described in Section 3.1. In the case of the PhysioNet2016 dataset, the two tasks match as it introduces a binary classification problem in the first place. For a fair comparison, all test data splits remain the same across each evaluation scheme. Finally, to assess the pretrained backbone encoder (and therefore the applied augmentations) on all available data, we repeat Steps 4 and 5 with the remaining datasets in a round-robin fashion (step 6 of the proposed framework).
@@
To thoroughly evaluate the proposed method we also repeat the entire experiment by leaving data out during Step 1 and pretraining the model with fewer data. This is also a means to concretely evaluate the proposed method on true OOD data, as they are completely hidden throughout the training process, either pretraining or downstream training. The best augmentation combinations are selected based on the effectiveness of the downstream models on the ‘In-Distribution’ test split. The same models are then evaluated on the hidden ‘OOD data‘.
3.5 Implementation Details

### v4

@@ II-C Model Robustness in 1D biosignal classification
Extracting generalized and robust representations has been one of the most important long-standing goals of machine learning [50]. To test and improve the generalizabilty and robustness of deep learning models, several fields have emerged, such as Domain Adaptation [51], Transfer Learning [52], and Domain Generalization [53, 4, 54], which use data from non-i.i.d. distributions or out-of-distribution (OOD) data for model evaluation. These distinct data distributions are referred to as data Domains, as they are most often the outcome of different data-generating processes. Here, we reference prior works which fall under the above research fields.
@@
Domain adaptation (DA) methods are the most popular choice when aiming to produce models with improved generalization capabilities in 1D biosignal classification. Similar to Transfer Learning (TL), DA algorithms leverage the representation learning capabilities of models pretrained on diverse source data distributions and fine-tune them on previously unseen but similar target distributions. The end-result is a model which remains robust under the distribution shift between source and target data. In healthcare, DA methods are proposed for a plethora of biosignal classification tasks. In [55], the authors propose pretraining vanilla CNN networks on a large corpus of raw ECG signals and demonstrate their effectiveness when fine-tuned on a smaller dataset for arrhythmia detection. In another more recent work, He et al. [56] propose tackling the distribution shift between different arrhythmia detection ECG datasets, by employing three distribution alignment mechanisms. To be more specific, they first extract spatio-temporal features from preprocessed signals, pass them through a graph convolutional neural network and finally align the resulting representations between source and target domains. Many studies have researched the advantages of DA and TL in EEG classification as well [57]. To name a few, the authors of [58, 59] and [60] propose aligning source and target distributions via adversarial learning for motor image EEG classification and epilepsy detection respectively, while [61] investigates several state-of-the-art DA algorithms for EEG-based emotion recognition.
@@
Finally, Domain Generalization (DG) algorithms aim to tackle model generalization at its core, as the main goal is to provide models that perform well across unseen datasets. Although limited, there has been a recent interest in biosignal DG research. In the field of PCG classification, the authors of [62] propose a ensemble classifier fusion method yields improved results in a DG setup for abnormal vs normal signal classification. To tackle the domain shift and variability present in ECG signals collected from different patients, [63] and [64] propose domain-adversarial model training for ECG and EEG classification, respectively. Most recently, the authors of [4] introduce a DG benchmark and a model which leverages representations from multiple layers of the network, inspired from [65], for 12-lead ECG and 64-channel EEG classification.
@@ III Materials and Methods
In this work, we introduce an evaluation framework for assessing the combination of various transformations, as well as the robustness of a network trained via contrastive SSL, against a fully supervised baseline model. This allows us to evaluate and identify the most effective signal transformations or augmentations within the PCG classification domain, while also highlight additional benefits of training an end-to-end classification model via Self-Supervised Contrastive Learning, such as improved model robustness and generalization.
@@
For the self-supervised training of all models, we adapt the SimCLR [66] methodology to 1D PCG representation learning and evaluate numerous augmentation combinations designed specifically for 1D signal processing tasks.
@@
In the following sections, we describe the datasets and preprocessing steps followed for all signals included in the study, provide a detailed description of the protocol for evaluating the significance of each combination of augmentations, as well as the generalization ability of a model pretrained with self-supervised contrastive learning. Finally, we present the architecture and relevant implementation details of the 1D CNN network implemented for PCG classification.
III-A Datasets & Downstream tasks
@@
The PCG signals used in this study originate from a total of 5 distinct public datasets and are available as “wav” files. These include the Ephnogram [67], FPCGDB [68], Pascal [69], PhysioNet2016 [14], and PhysioNet2022 [15] databases. Here, we describe each dataset separately and provide all necessary details regarding the preprocessing steps followed to homogenize all signals into a common format.
@@
The Ephnogram [67] database contains a total of 69 multi-channel, simultaneous ECG and PCG recordings from 24 healthy adults aged between 23 and 29, acquired during physical activity and rest conditions. Most recordings (62 during exercise) are 30min long, while the duration of the remaining 7 (rest) is about 30 seconds. As specified by the researchers, the dataset includes poor sample signals making it appropriate for developing noise-resilient methods. As this dataset was proposed and gathered for multi-modal signal processing purposes and not for pathology detection or classification, it does not include labeled samples. For our study, we use the unlabeled PCG signals for the SSL pretraining of our models. All recordings were sampled at 8 kHz.
@@
The Fetal PCG Database (FPCGDB) [68], as its name suggests, is a dataset comprising 26 20-minute fetal phonocardiographic signals collected from healthy women during the final months of their pregnancies. The signals were sampled at 333 Hz. Similar to the Ephnogram dataset, these recordings were gathered for the development of fetal PCG signal simulation and fetal heart rate extraction algorithms, rather than for classification purposes, and thus do not include labeled samples. Additionally, many recordings are affected by noise, originating either from the pregnant body, such as muscular movements or placental blood turbulence, or from the surrounding environment.
@@
The Pascal [69] Heart Sound Classification Challenge was introduced for developing algorithms for the automatic screening of cardiac pathologies. The challenges contains two tasks; one for heart sound segmentation and one for classification. In our work, we use the publicly available labeled data for model training and evaluation which was gathered in clinical and in-the-wild settings. Notably, the 832 PCG signals are classified as either Normal, Murmur, Extra Heart Sound, Artifact or Extrasystole [^2] Each different category is described in the official challenge description (https://istethoscope.peterjbentley.com/heartchallenge/index.html).. The recording lengths vary from 1 to 30 seconds and are sampled at 195 Hz. As each sample is annotated with a single label, the above task is a multi-class classification problem.
@@
The remaining two datasets were also introduced as publicly available challenges for PCG classification. Similarly to Pascal, the objective of the PhysioNet2016 [14] Challenge aims at urging researchers to develop robust algorithms for differentiating between normal and abnormal heart sound recordings as a binary classification problem. The data included for model training in the dataset contains a total of 3,126 PCG recordings collected from a single precordial location from both healthy subjects and patients suffering from a variety of heart defections, lasting from 5 to around 120 seconds. Once again, the recordings were gathered in both clinical and non-clinical environments, with a sample rate of 2 kHz.
@@
Finally, the PhysioNet2022 Challenge [15] was designed for the development of algorithms that can detect heart murmur in PCG signals. Building on the 2016 challenge, the 2022 edition contains a diverse dataset of heart sound recordings collected from multiple auscultation locations. The dataset includes 3,163 recordings obtained from 942 patients, with each patient contributing one or more PCG signals recorded from one of the four heart valves or another prominent auscultation location. Each recording is classified into one of three categories: present (heart murmur detected), absent (no heart murmur detected), or unknown (e.g., due to noise or corrupted audio), rendering this a multi-class classification task. All recordings were sampled at a rate of 4 kHz.
@@
Each dataset contains signals of varying length, sampling rate and label classes. To accommodate model training in our framework, we follow a similar process to [27] for converting all signals into a common format. As a first step and to avoid transient noise artifacts, we discard the first and final 2 seconds of each PCG recording. The resulting signal is split into 5 second overlapping windows with a 50% overlap. After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether or not it derives from a normal heart sound recording [^3] In this study, all recordings initially labeled as unknown are considered abnormal.. For the SSL pretraining stage we assign pseudolabels to each window, as described in Section III-B. The assignment of binary labels to each window enables the evaluation of trained models across datasets, which do not necessarily share similar sample classes. For example, a classifier trained to detect normal PCG signals in the Pascal dataset, can also be evaluated without further training on samples from the PhysioNet2022 dataset. The above evaluation process is described in detail in Section III-D.
III-B Self-Supervised Contrastive Learning for PCG Representation Learning
@@
The main goal of this study is to first of all verify the use of SSL contrastive learning, but also evaluate the importance of augmentation selection for robust PCG representation learning. The proposed framework consists of two separate tasks; a Pretraining pretext task and a Downstream task, as depicted in Figure 1.
@@
Assume a 1D PCG signal $x\in\mathbb{R}^{1\times L}$ sampled from $\mathcal{X}$, where $L$ is the length of the signal. The aim of the contrastive SSL pretext task is to learn robust representations of $x$ from unlabeled PCG signals, by maximizing the agreement between two different augmented versions of itself (positive pair) and simultaneously minimizing the agreement between entirely separate augmented windows. Essentially, a model is trained to predict whether two augmented signals are the result of different transformations applied on the original sample. Formally, let $t(\cdot)$ be a transformation or augmentation function drawn from a family of functions $\mathcal{T}$. During the pretraining phase, an audio window $x$ is augmented to two different versions $\tilde{x}$ and $\tilde{x}^{\prime}$, after the application of two transformation functions $t\sim T$ and $t^{\prime}\sim T^{\prime}$ respectively. Therefore, an initial batch of $N$ sample windows $(\tilde{x}_{1},\tilde{x}_{2},\dotsc,\tilde{x}_{n})$ results into an augmented batch of $2N$ signals $(\tilde{x}_{1},\tilde{x}_{2},\dotsc,\tilde{x}_{n},\tilde{x}_{1}^{\prime},\tilde{x}_{2}^{\prime},\dotsc,\tilde{x}_{n}^{\prime})$, where $\tilde{x}_{i}=t(\tilde{x})$ and $\tilde{x}_{i}^{\prime}=t^{\prime}(\tilde{x}_{i})$ for $i\in[1,N]$. As mentioned, each augmented pair of signals resulting from the same original window is considered a positive pair, e.g $(\tilde{x}_{1},\tilde{x}_{1}^{\prime})$, while pairs of different original windows comprise negative pairs, e.g $(\tilde{x}_{1},\tilde{x}_{2}^{\prime})$. Following its augmentation, the $2N$ signal batch is passed through a neural network encoder model $E(\cdot)$ (Figure 2), where each signal is encoded into a representation $\tilde{h}=E(\tilde{x})$. Subsequently, the resulting representations are flattened and passed through a fully connected linear layer $G(\cdot)$ and projected into 128-dimensional vectors $\tilde{z}=G(\tilde{h})$, as proposed in [70].
@@
For the contrastive prediction task, we employ the NT-Xent loss function introduced in [30] and implemented in the SimCLR framework [70]. Specifically, given a set $\{\tilde{x}_{k}\}$ which includes an augmented positive pair $\tilde{x}_{i}$ and $\tilde{x}_{j}$, a model trained via contrastive SSL aims to identify $\tilde{x}_{j}$ in $\{\tilde{x}_{k}\}_{k}\neq i$ for a given $\tilde{x}_{i}$. The contrastive loss for a positive pair $(\tilde{x}_{i},\tilde{x}_{j})$ is formulated as follows:
@@
III-C Augmentations for PCG Representation Learning
@@ III-D Evaluation Protocol
In the first step of the framework, we create a pool of 5 sec unlabeled audio windows from all available datasets, namely Pascal, PhysioNet2016, PhysioNet2022, Ephnogram and Fpcgdb. After the preprocessing step has been completed and all signals are homogenized in the common format described in Section III-A, they are randomly shuffled and split into training and validation splits (80% and 20% respectively). In the second step, the split signals are cut into batches. Each sample in the batch is then duplicated and augmented by applying either one or a combination of the functions described in Section III-C. The augmented samples are then passed through a Neural Network encoder, which is trained via the contrastive pretraining task (Step 3). After the SSL backbone model has been trained, its weights are frozen.
@@
In Step 4, a classification head with learnable weights is attached to the pretrained encoder which is then trained on data from one of the labeled datasets (Pascal, PhysioNet2016 or PhysioNet2022). Specifically, each of these datasets is split into training, validation and test splits with a standard 70%-20%-10% ratio and the trained model is evaluated on the test split only. Since each model is evaluated on test data originating from the same distribution as its training data, we call this the ”In-Distribution” evaluation. Inspired by recent work in biosignal domain generalization [4], we also design an evaluation protocol for assessing the robustness of a trained classifier on out-of-distribution data, which are drawn from entirely different datasets. Since the data has been completely left-out during downstream training, the above process is followed for evaluating the generalization ability of a model. Specifically, we now also freeze the weights of the attached classification layers and test the model on data drawn from the remaining 2 labeled datasets. For example, if the network is trained for the Pascal downstream task, it is evaluated on both in-distribution data (Pascal) and on out-of-distribution data (PhysioNet2016 and PhysioNet2022). We call this step the out-of-distribution or OOD evaluation step. However, in order to evaluate the same model on each separate dataset, all data must share common labels. To that end, we implement two different downstream task types: ‘All’ and ‘Binary’. In the ‘All’ downstream task, the original labels of each dataset are assigned to each audio window, while in the ‘Binary’ task a signal is labeled as either ‘normal’ or ‘abnormal’, as described in Section III-A. In the case of the PhysioNet2016 dataset, the two tasks match as it introduces a binary classification problem in the first place. For a fair comparison, all test data splits remain the same across each evaluation scheme. Finally, to assess the pretrained backbone encoder (and therefore the applied augmentations) on all available data, we repeat Steps 4 and 5 with the remaining datasets in a round-robin fashion (step 6 of the proposed framework). The significance of each augmentation and transformation is evaluated based on: (a) their occurence frequency in the top performing models in both In-Distribution and OOD evaluation splits, and (b) their effect on model training, as calculated by Cohen’s d measure [71]. Additional information and details regarding augmentation evaluation are provided in Section IV-A.
@@
To thoroughly evaluate the proposed method we also repeat the entire experiment by leaving data out during Step 1 and pretraining the model with fewer data. This is also a means to concretely evaluate the proposed method on true OOD data, as they are completely hidden throughout the training process, either pretraining or downstream training. The best augmentation combinations are selected based on the effectiveness of the downstream models on the ‘In-Distribution’ test split. The same models are then evaluated on the hidden ‘OOD data’.
III-E Implementation Details